Circular No. 8833

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2007 H3 (GARRADD) G. J. Garradd reports his discovery of a comet on CCD images obtained

with the 0.5-m Uppsala Schmidt at Siding Spring (discovery observation tabulated below), the object showing a faint tail 15′′ long in p.a. 10^{o} on the 30-s survey exposure. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, N. Teamo and S. Hönig report that their CCD frames taken with a 0.41-m f/8 reflector at Punaauia, Tahiti, on Apr. 23.3 UT show a 4′′ coma and an 8′′ tail in p.a. ≈ 35^{o}.

2007 UT  α_{2000}  δ_{2000}  Mag.
Apr. 22.51406  11^{h}50^{m}44.^{s}52  −42^{o}53′06′′.8  17.4

The available astrometry (including prediscovery Siding Spring observations from Mar. 28), the following preliminary elliptical orbital elements, and an ephemeris appear on MPEC 2007-H55.

T = 2007 Aug. 15.509 TT  ω = 349.^{o}968
e = 0.47547  Ω = 263.731 }2000.0
q = 1.83355 AU  i = 25.225
a = 3.49558 AU n^{o} = 0.150808 P = 6.54 years

SUPERNOVA 2006jc IN UGC 4904 T. Minezaki, Y. Yoshii, and K. Nomoto, University of Tokyo, report

the following near-infrared photometric magnitudes of SN 2006jc (cf. IAUC

8762; http://www.astronomerstelegram.org/?read=961) using the mul-

ticolor imaging photometer mounted on the MAGNUM 2-m telescope: Apr. 9.43 UT, K = 16.2 ± 0.1; 10.40, H = 17.8 ± 0.1; 10.43, J = 19.5 ± 0.3. Limiting optical magnitudes 3σ upper limit: Apr. 9.43, V = [22.5; 10.40, R = [22.5; 10.43, I = [21.7.

K. S. Kawabata, Hiroshima University; K. Maeda, Max Planck Institute for Astrophysics; M. Tanaka, N. Tominaga, and K. Nomoto, University of Tokyo; and T. Hattori, National Astronomical Observatory of Japan, report that optical imaging and PSF photometry of the peculiar type-Ib supernova 2006jc was performed with the Subaru 8.2-m telescope (+ FO-CAS) on Apr. 10.3 UT, yielding magnitudes B = 24.0 ± 0.1, V = 23.4 ± 0.2, and R = 23.3 ± 0.2.

2007 April 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green